OpenStudy (anonymous):

(x-1)^7/2 - (x-1)^3/2 factor and simplify. can u explain how to do it

OpenStudy (anonymous):

holy cow...

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

(x-1)^3/2((x-1)^2-1) =(x+1)(x-1)^5/2

OpenStudy (anonymous):

not the right answer

OpenStudy (anonymous):

oops! calculation error! it is (x-1)^3/2(x-2)x

OpenStudy (anonymous):

yes thats the right answer. my teacher gave us a packet with the problems and answers already. i just need to knw how to work it step by step.

OpenStudy (anonymous):

can u show me how u got that

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

From: (x-1)^(7/2) - (x-1)^(3/2) we can factor out (x-1)^(3/2), so we get: (x-1)^(3/2) ((x-1)^2 - 1) (if you don't see why we reach this step let me know) Then, working out the right hand portion of this we get (x-1)^(3/2) (x^2 - 2x + 1 - 1) The ones cancel (x-1)^(3/2) (x^2 - 2x) Factor out an x from the right. (x-1)^(3/2) (x) (x - 2)

OpenStudy (anonymous):

i get the fist step, how to factor it out. the second step i dont get how u got ((x-1)^2-1

OpenStudy (anonymous):

Okay, we have the equation: \[(x-1)^{\frac{7}{2}} - (x-1)^{\frac{3}{2}}\] when we factor out the \[(x-1)^{\frac{3}{2}}\] what we leave behind needs to be whatever you would multiply the bit we factored out by to obtain the original equation. ie: \[(x-1)^{\frac{3}{2}} (???) = (x-1)^{\frac{7}{2}} - (x-1)^{\frac{3}{2}}\] okay, so we need to figure out what the ??? is. To get the first term of the original equation, (x-1)^(7/2), what would we need to multiply (x-1)^(3/2) by? ie.. \[(x-1)^{\frac{3}{2}} \times a = (x-1)^{\frac{7}{2}}\] what is a? If you have trouble seeing what a should be, think about the following: \[x^{3} \times a = x^{7}\] what is a here? It should be clear that \[a = x^{4}\] Because when we multiply powers with like bases, we add the exponents. so a in the first equation is: \[a = (x-1)^{2}\] since (3/2) + 2 = (3/2) + (4/2) = (7/2) So we know the first part of ??? is (x-1)^2, we need another part though, to get the -(x-1)^(3/2) from the orginal equation. This bit is simple, since we just need to multiply our factor by -1 to get this. So... \[(x-1)^{\frac{3}{2}} ((x-1)^{2} - 1) = (x-1)^{\frac{7}{2}} - (x-1)^{\frac{3}{2}}\]